# METAL: A Multi-Agent Framework for Chart Generation with Test-Time Scaling

### v1

@@ Abstract
Chart generation aims to generate code to produce charts satisfying the desired visual properties, e.g., texts, layout, color, and type. It has great potential to empower the automatic professional report generation in financial analysis, research presentation, education, and healthcare. In this work, we build a vision-language model (VLM) based multi-agent framework for effective automatic chart generation. Generating high-quality charts requires both strong visual design skills and precise coding capabilities that embed the desired visual properties into code. Such a complex multi-modal reasoning process is difficult for direct prompting of VLMs. To resolve these challenges, we propose METAL (Multi-agEnT frAmework with vision Language models for chart generation), a multi-agent framework that decomposes the task of chart generation into the iterative collaboration among specialized agents. METAL achieves 5.2% improvement in accuracy over the current best result in the chart generation task. The METAL framework exhibits the phenomenon of test-time scaling: its performance increases monotonically as the logarithmic computational budget grows from 512 to 8192 tokens. In addition, we find that separating different modalities during the critique process of METAL boosts the self-correction capability of VLMs in the multimodal context.
@@ 1 Introduction
In this paper, we present METAL, a multi-agent framework designed for chart generation. Our framework decomposes this complex multimodal reasoning task into four specialized roles, each handled by a specialized agent: (1) Generation Agent: Responsible for the initial translation of chart images into the corresponding code. (2) Visual Critique Agent: Analyzes and identifies visual differences between the reference chart and the generated output. (3) Code Critique Agent: Reviews the generated code and suggests improvements to better match the reference chart. (4) Revision Agent: Implements code modifications based on the combined feedback from both critique agents. During inference, these agents collaborate iteratively, critiquing and refining the code until the rendered chart achieves the desired accuracy.
As illustrated in Figure 1, current state-of-the-art VLMs, such as GPT-4o, often fail to accurately interpret and reproduce the intricate visual elements and relationships embedded in reference charts. Existing solutions, such as Best-of-N and Hint-enhanced Wang et al. (2024), have not effectively improved upon direct prompting of VLMs. The core challenge in leveraging VLMs for chart generation lies in effectively integrating visual comprehension with code synthesis. This complex task exceeds the capabilities of the single model or single agent. In contrast to existing methods, our approach delivers more concrete and targeted feedback, and iteratively refines outputs through the multi-agentic framework, leading to enhanced chart generation performance.
Experiment results show that our framework improves chart generation accuracy by over 11.33%, demonstrating its potential to significantly enhance VLMs’ ability to integrate visual understanding with code synthesis.
Furthermore, we have two key findings through in-depth analysis: (1) Test-time scaling in the METAL framework: We found that there is a near-linear relationship between the performance and the logarithm of the computational budget in experiments. Specifically, the performance of METAL increases monotonically as the logarithm of the computational budget grows from 512 to 8192 tokens. (2) Modality-tailored critiques enhance self-correction: We observe that explicitly separating different modalities during the critique process—such as visual evaluation and code analysis—substantially enhances the multimodal self-correction capabilities of VLMs. Ablation study shows that METAL with the separate-critique design achieves a 5.16% improvement over the single-critique baseline.
@@ 3.2 METAL
We design a heuristic-based verification function to evaluate the chart quality. Let $m_{j}$ be the verification metrics for $j=1,2,3$, and let $\theta^{t}$ be dynamic thresholds. Then,
@@
More details on the implementation of the verifier, such as each verification metric, are introduced in Appendix A.2.
@@ 3.3 Inference Procedure
During inference, METAL iteratively refines the generated code until the rendered chart meets a predefined quality threshold. The refinement process is as follows:

### v2

@@ Abstract
Chart generation aims to generate code to produce charts satisfying the desired visual properties, e.g., texts, layout, color, and type. It has great potential to empower the automatic professional report generation in financial analysis, research presentation, education, and healthcare. In this work, we build a vision-language model (VLM) based multi-agent framework for effective automatic chart generation. Generating high-quality charts requires both strong visual design skills and precise coding capabilities that embed the desired visual properties into code. Such a complex multi-modal reasoning process is difficult for direct prompting of VLMs. To resolve these challenges, we propose METAL (Multi-agEnT frAmework with vision Language models for chart generation), a multi-agent framework that decomposes the task of chart generation into the iterative collaboration among specialized agents. METAL achieves a 5.2% improvement in F1 score over the current best result in the chart generation task. Additionally, METAL improves chart generation performance by 11.33% over Direct Prompting with LLaMA 3.2-11b. Furthermore, the METAL framework exhibits the phenomenon of test-time scaling: its performance increases monotonically as the logarithm of computational budget grows from $2^{9}$ to $2^{13}$ tokens.
@@ 1 Introduction
As illustrated in Figure 1, current state-of-the-art VLMs, such as GPT-4o, often fail to accurately interpret and reproduce the intricate visual elements and relationships embedded in reference charts. Existing solutions, such as Best-of-N and Hint-enhanced Wang et al. (2024), have not effectively improved upon direct prompting of VLMs. The core challenge in leveraging VLMs for chart generation lies in effectively integrating visual comprehension with code synthesis. This complex task exceeds the capabilities of the single model or single agent.
In this paper, we present METAL, a multi-agent framework designed for chart generation. Our framework decomposes this complex multimodal reasoning task into four specialized roles, each handled by a specialized agent: (1) Generation Agent: Responsible for the initial translation of chart images into the corresponding code. (2) Visual Critique Agent: Analyzes and identifies visual differences between the reference chart and the generated output. (3) Code Critique Agent: Reviews the generated code and suggests improvements to better match the reference chart. (4) Revision Agent: Implements code modifications based on the combined feedback from both critique agents. During inference, these agents collaborate iteratively, critiquing and refining the code until the rendered chart achieves the desired quality.
In contrast to existing methods, our approach delivers more concrete and targeted feedback, and iteratively refines outputs through the multi-agentic framework, leading to enhanced chart generation performance. Experiment results show that our framework achieves a 5.2% improvement over the current best result in the chart generation task. Additionally, METAL improves chart generation performance by 11.33% over Direct Prompting with LLaMA 3.2-11b, demonstrating its potential to significantly enhance VLMs’ ability to integrate visual understanding with code synthesis.
Furthermore, we have two key findings through in-depth analysis: (1) Test-time scaling in the METAL framework: Recent works suggest the potential of scaling the number of tokens (computational budget) to enhance the reasoning performance of LLMs (Snell et al., 2024; Muennighoff et al., 2025). Specifically, utilizing more tokens during inference may lead to improved performance. In this work, we found that there is a near-linear relationship between the performance and the logarithm of the computational budget in experiments. Specifically, the performance of METAL increases monotonically as the logarithm of the computational budget grows from $2^{9}$ to $2^{13}$ tokens. (2) Modality-tailored critiques enhance self-correction: We observe that explicitly separating different modalities during the critique process—such as visual evaluation and code analysis—substantially enhances the multimodal self-correction capabilities of VLMs. Ablation study shows that METAL with the separate-critique design achieves a 5.16% improvement over the single-critique baseline.
@@ 3.2 METAL
We design a heuristic-based verifier to evaluate the chart quality. The goal of verifier is to provided external source of feedback to guide four agents collaborate more efficiently.
Let $m_{j}$ be the verification metrics for $j=1,2,3$, and let $\theta^{t}$ be dynamic thresholds. Then,
@@
Here, we developed three verification metrics—color ($m_{1}$), text ($m_{2}$), and overall structure ($m_{3}$)—to quantify the similarity between the reference and generated charts. If the generated chart meets the desired quality (i.e., the verification score of each verification metrics ($m_{1}$, $m_{2}$, $m_{3}$) exceeds the predefined threshold), the system triggers an early stop. More details on the implementation of the verifier, such as each verification metric, are introduced in Appendix A.2.
Algorithm 1 Inference Procedure of METAL
1:$y_{0}\leftarrow G(x_{\text{ref}})$triangleright Generate initial code
2:while $t<T_{\max}$ do
3:     $O(y_{t})\leftarrow$ Render chart from $y_{t}$
4:     $v_{t}\leftarrow V(O(y_{t}),x_{\text{ref}})$triangleright Visual critique
5:     $c_{t}\leftarrow C(y_{t})$triangleright Code critique
6:     if $\forall j:\;m_{j}(O(y_{t}),x_{\text{ref}})\geq\theta^{t}$ then
7:         break $\triangleright$ Verification passed
8:     else
9:         $y_{t+1}\leftarrow R(y_{t},v_{t},c_{t})$triangleright Revise code
10:     end if
11:     $t\leftarrow t+1$
12:end while
13:return $y_{t}$
@@ 3.3 Inference Procedure
Algorithm 1 illustrates the inference procedure of METAL. During inference, METAL iteratively refines the generated code until the rendered chart meets the predefined verification threshold or reaches the maximum attempts limit $T_{\max}$. The refinement process is as follows: